Ticket Number: HEA-NOC/20031124-7 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20031124 18:14 UTC Problem Start: 20031127 10:00 UTC
Ticket Closed: Problem End: 20031127 12:00 UTC
On Thursday, 27/11/03, there will be a short outage on ftp.heanet.ie to
facilitate a physical move to a new cabinet, a kernel upgrade, and the
migration of the storage array to an NFS backend server. This outage is
expected to last in the region of 30 minutes within the maintainence
window of 2 hours.
All access to ftp.heanet.ie will be limited during the outage.
20031124 18:06 CMC Opening ticket.
20031127 10:00 CMC ftp.heanet.ie powered down
20031127 10:40 CMC ftp.heanet.ie powered back up
20031129 21:45 CMC due to NFS scalability problems, ftp.heanet.ie has been
migrated to backend server.
This ticket can be monitored at http://www.hea.net/tickets/20031124-7